ON THE HOMOLOGY COBORDISM GROUP OF HOMOLOGY 3-SPHERES

NIKOLAI SAVELIEV

Abstract. In this paper we present our results on the homology cobordism group Θ_{ℤ}^{3} of the oriented integral homology 3-spheres. We specially emphasize the role played in the subject by the gauge theory including Floer homology and invariants by Donaldson and Seiberg – Witten.